The side lengths of the triangle are increased by 100%. The height is 3cm, the base is 4cm, and the hyptoenuse is 5cm. What is the percent of increase in area from the original triangle to the new triangle?

Answer by checkley79(3341)

AREA=BH/2
AREA=3*4/2
AREA=12/2=6 cm.
AREA=6*8/2=48/2=24 cm.
24-6=18 cm. ANS. FOR THE INCREASE.